OpenStudy (anonymous):

Given one endpoint X (-2,7) AND M (3,7) Find Ythe other endpoint. Pleaseee helpp :>

OpenStudy (campbell_st):

M is a midpoint then 3 = (-2 + x)/2 7 =(7 + y)/2 solve for x and y 6 = -2 + x 14 = 7 + y I'll let you finish

OpenStudy (anonymous):

I dont see what to do next @campbell_st

OpenStudy (campbell_st):

6 = -2 + x 14 = 7 + y add 2 to both sides subtract 7 from both sides 8 = x 7 = y the endpoint is (8, 7)

OpenStudy (anonymous):

If thats your answer , it isnt correct . not one of the answer choices .

OpenStudy (campbell_st):

Is M the midpoint..?

OpenStudy (anonymous):

i dont think so , i think its just a letter given .

OpenStudy (anonymous):

@Limitless heyy . do you think you know his mistake in this one?

OpenStudy (campbell_st):

If M isn't a midpoint... and you have no other information about M... you can't find the point Y

OpenStudy (anonymous):

I believe \((3,7)\) has to be the midpoint. Otherwise there is not a sensible question here.

OpenStudy (anonymous):

this is a practice question given from the school

OpenStudy (anonymous):

\[\text{mid}(A,B)=\left(\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2}\right)\] So, \[(3,7)=\left(\frac{-2+x_2}{2}, \frac{7+y_2}{2}\right)\] Can you solve from here?

OpenStudy (anonymous):

@campbell_st is correct, it seems.

OpenStudy (anonymous):

the answer choice tht i was going with was (-8,-1) ?

OpenStudy (campbell_st):

what solutions were you given...?